# Massively multiplexed nanoscale magnetometry with diamond quantum sensors

### v1

@@
Kai-Hung Cheng Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Zeeshawn Kazi Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Jared Rovny Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Bichen Zhang Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Lila Nassar Princeton University, Department of Physics, Princeton, NJ 08544, USA    Jeff D. Thompson Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Nathalie P. de Leon Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA
(August 21, 2024)
@@
Single nitrogen vacancy (NV) centers in diamond have been used extensively for high-sensitivity nanoscale sensing, but conventional approaches use confocal microscopy to measure individual centers sequentially [38, 27, 22, 21, 20, 13], limiting throughput and access to non-local physical properties. Here we design and implement a multiplexed NV sensing platform that allows us to read out many single NV centers simultaneously using a low-noise camera. Using this platform, we coherently manipulate and read out the spin states of hundreds of individual NV centers in parallel, achieving comparable magnetic field sensitivity to confocal measurements. We also implement a parallelized version of spin-to-charge-conversion readout for low NV center spin state readout noise [36] and use it to demonstrate multiplexed covariance magnetometry [34], in which we measure six two-point magnetic field correlators from four NV centers simultaneously. The number of correlators we can measure is limited only by the available laser power, opening the door to massively multiplexed covariance magnetometry. Our platform significantly increases the throughput and broadens the applications of nanoscale sensing using diamond quantum sensors [32, 4].
@@
Nitrogen vacancy (NV) centers in diamond have long spin coherence times across a wide range of temperatures combined with optical spin initialization and readout, enabling broad applications in nanoscale sensing of quantities such as magnetic field, electric field, and temperature [38, 27, 22, 21, 20, 13]. NV centers are typically deployed either as individual sensors in bulk diamond [38, 27, 22, 37, 20, 10], scanning tip geometries [26, 2], or in dense ensembles of NV centers that are not individually resolvable [16, 15, 6, 17, 31]. Resolvable NV centers offer nanometer spatial precision [30] and high sensitivity even for NV centers within a few nanometers of the surface with proper surface preparation [35, 40] but suffer from slow measurements due to the high readout noise of the NV center spin state [12, 5]. NV center ensembles offer fast, wide-field measurements by averaging over the signal measured by many NV centers [16, 15, 5, 18] at the cost of micron-scale spatial resolution imposed by the diffraction limit.
@@
In this work, we demonstrate a multiplexed NV sensing platform that combines the best of these modalities, realizing parallel, wide-field measurements of many individually resolvable NV centers using a low-noise camera. We first demonstrate parallel continuous-wave (CW) ODMR of hundreds of resolvable NV centers, achieving a magnetic field sensitivity within a factor of 5 of confocal measurements of single centers in the same sample. We extend this technique to spatially resolved spin measurements, performing parallel relaxometry ($T_{1}$ measurements) of 70 NV centers. Next, we demonstrate enhanced parallel spin readout by implementing spin-to-charge conversion (SCC) using a spatial light modulator (SLM) to focus a high-power ionization laser onto multiple NV centers simultaneously. We realize efficient ionization and multiplexed single-shot charge state readout with a fidelity of 88%, which enables multiplexed SCC spin readout with low readout noise, around ten times the quantum projection noise limit. This architecture allows us to measure magnetic field correlations between four individually resolved NV centers in parallel. The scalability of the multiplexed SCC is limited by the ionization laser power, and we expect that these techniques can be extended to hundreds of NV centers in future work, enabling parallel nanoscale sensing of non-local properties at many length scales [32].
@@
Our multiplexing platform (Fig. 1a) uses an electron-multiplying charge-coupled device (EMCCD) camera to detect NV center fluorescence and a microwave printed-circuit board in front of the objective to deliver microwave excitation [41]. Using CW wide-field green (532 nm) excitation, we image several hundred NV centers with approximately 2 seconds of integration time on the camera (Fig. 1b). We demonstrate parallel spin control and readout by measuring the optically-detected magnetic resonance (ODMR) spectrum of all the NV centers in the field of view. We simultaneously measure several hundred ODMR spectra in parallel (Fig. 1d) by recording the fluorescence emitted by each NV center under CW excitation as a function of applied microwave frequency and normalizing by a reference measurement with no applied microwaves (Fig. 2a). NV centers exist in four possible crystallographic orientations [38], and three distinct orientations are observed here (Fig. 1d, different colors) because two are degenerate under the static magnetic field alignment. From the ODMR spectra, we calculate the DC magnetic field sensitivity of each individual NV center and find the median DC sensitivity is 24 µT Hz-1/2 (Fig. 2b) compared to a median sensitivity of 5.5 µT Hz-1/2 obtained with confocal measurements of the same diamond sample (Extended Data Figs. 1 and 2).
@@
Parallel readout also enables the multiplexed measurement of NV center dynamics. Pulsed measurements are challenging because the EMCCD camera data transfer time of 10 ms is much longer than the typical readout time for NV center spin contrast of 300 ns, determined by the lifetime of the shelving state [12]. For wide-field pulsed measurements using the camera, we repeat the experiment multiple times while integrating the fluorescence on the camera (Fig. 2d[18]. This reduces the overhead of frame transfer time and suppresses the noise from camera readout but lowers the fluorescence contrast between the NV spin states because the camera also records fluorescence during the spin reinitialization between experiments. With this approach, we are able to measure Rabi oscillations of many NV centers simultaneously (Fig. 2c). We demonstrate the application of parallel pulsed experiments to measure the spin relaxation time $T_{1}$ of over 70 NV centers in parallel (Fig. 2e). The measured $T_{1}$ times in this sample, which has NV centers implanted within approximately 10 nm of the surface, range from a few hundred microseconds to over 4 ms, consistent with past measurements of similarly prepared samples [35, 8].
@@
We next demonstrate multiplexed low noise readout of the NV center spin state using SCC, which is critical for covariance magnetometry [34] as the time required to measure a magnetic field correlation signal with a given signal-to-noise ratio scales as $\sigma_{\text{R}}^{4}$ where $\sigma_{\text{R}}$ is defined as the single-shot spin state readout noise [5]. SCC achieves lower $\sigma_{\text{R}}$ compared to green readout by using a high-power laser pulse to selectively ionize the NV center based on its spin state, followed by a high-fidelity charge state readout [36]. The selective ionization step typically requires $\sim 10$ mW of optical power per NV center, so implementing SCC with a wide-field laser beam would require $>$100 W of power for a 20 µm $\times$ 20 µm region. Thus, to realize multiplexed SCC in a power efficient manner, we use an SLM to create a programmable pattern of laser beams that only excite targeted NV centers (Fig. 1c). For simplicity, we use the same illumination pattern and wavelength (orange, 594 nm) for both ionization and charge state readout.
@@
We implement multiplexed charge state readout with the SLM and the camera through the following steps, outlined in Fig. 3a-c (Methods Section I). First, we take an image of all the NV centers using wide-field green excitation and identify the NV centers that we want to target (Fig. 3a). Second, we run a weighted Gerchberg-Saxton algorithm to generate a phase pattern on the SLM to create the laser spots at the fitted NV center locations (Fig. 3b[29]. To read out the charge state, we excite the NV centers with a low-power orange laser and measure the photon counts per exposure by thresholding individual pixels in a region around each NV center and then summing the thresholded counts (Fig. 3c). In this low-photon regime, we binarize the image such that each pixel measures either 0 or 1 photons and do not account for higher per-pixel counts, which allows us to take full advantage of the camera’s readout statistics [7]. A representative histogram of photon counts per camera frame is shown in Fig. 3d (Extended Data Fig. 3), revealing two peaks that correspond to the NV0 (dark) and NV- (bright) charge states under orange illumination. The measured charge state readout fidelity using the camera (88.3%) is similar to that obtained using confocal readout of the same sample (86.3%) with the same objective lens and readout parameters (Extended Data Fig. 4).
@@
Having achieved multiplexed charge-state measurements using targeted excitation, we then add a spin-selective ionization pulse to implement multiplexed SCC by modulating the power of the orange laser between ionization ($\sim$10 mW per spot) and readout ($\sim$20 µW per spot). We compare the charge state readout statistics for a representative NV center spin prepared in either the $m_{s}=0$ or $m_{s}=-1$ spin state to benchmark $\sigma_{\text{R}}$ (Fig. 3e). The spin readout noise for the data shown in Fig. 3e is $\sigma_{\text{R}}\approx 12$, a four-fold improvement over green readout in a confocal microscope [5] and a ten-fold improvement over wide-field green readout (Extended Data Fig. 5). However, our scheme performs worse than previous implementations of SCC using a red (637 nm) laser for ionization ($\sigma_{\text{R}}\sim 5$[34], which we attribute to a higher charge state recombination rate from the orange laser [3], in agreement with a rate equation model [39] (Extended Data Fig. 6).
@@
Equipped with multiplexed SCC, we demonstrate parallelized covariance magnetometry by reconstructing non-local magnetic field correlators $\braket{B(\vec{x}_{i})B(\vec{x}_{j})}$ from simultaneous measurements of NV centers at different locations $\vec{x}_{i}$ [34]. We target 4 NV centers (Fig. 4a) from which we simultaneously measure (${4\choose 2}=6$ two-point correlators. To demonstrate that the measured correlations result from spin dynamics rather than spurious correlations [34], we use a separate green laser to prepare one of the NV centers (NV4) in the opposite spin state from the other three (NV1-3) (Fig. 4a,b). To verify our ability to detect two-point correlators with our achievable readout noise, we first measure correlations introduced by direct NV spin driving rather than noise sensing (Fig. 4b). This allows us to rapidly drive the NV centers to mitigate the effect of decoherence under free evolution, as well as to circumvent quantum projection noise as we are able to intentionally drive each NV center between pure states $m_{s}=0$ and $m_{s}=-1$. In this driven-spin correlation experiment, we interleave two pulse sequences: in the even-numbered experiments, we simultaneously drive all the NV centers around the Bloch sphere by an angle $\theta$, while in the odd-numbered experiments we drive the NV centers by an angle $\theta+\pi$ (Fig. 4b). At the end of each experiment we read out a single camera frame, simultaneously measuring the four signals $s_{i}$ where $i=1...4$ indexes the NV center. Lastly, we use the set of acquired signals across experiments $S_{i}=\{s_{i,k}\}$, where $k$ indexes the experiment number, to compute the Pearson correlation for each NV pair: $r_{ij}=\text{Cov}(S_{i},S_{j})/(\sigma_{i}\sigma_{j})$, where Cov is the covariance and $\sigma_{i}$ is the standard deviation of $S_{i}$. From the computed Pearson correlations, we subtract off a uniform background correlation measured using an independent calibration experiment (Extended Data Fig. 7); the resulting correlations are shown in Fig. 4c. As expected for each pair of NV centers prepared in the same initial spin state (NV1-3), the correlations between the measured signals are maximized when $\theta=n\pi$ for integer $n$, as the two spins flip synchronously between $m_{s}$ = 00 and $-1$, and are minimized for $\theta=(n+1/2)\pi$, since the readout projects the two spins randomly and independently onto $m_{s}$ = 00 or $-1$. Conversely, each NV center’s correlation with NV4 is negative, as NV4 has been prepared in the opposite orientation and thus acquires an opposite phase. In addition, the amplitude of the correlation oscillation indicates the maximum measurable correlation, which is $1/\sigma_{R_{i}}\sigma_{R_{j}}$. From Fig. 4c, we estimate $\sigma_{R_{i}}=$ $15\pm 1$, consistent with the data in Fig. 3f.
@@
We also use multiplexed covariance magnetometry to perform covariance noise spectroscopy of an applied AC magnetic field [34]. As before, we prepare NV1-3 in the same spin state and prepare NV4 in the opposite state. We then use an XY8 sensing sequence [28] with a fixed interpulse spacing, and sweep the frequency of the applied AC signal (Fig. 4d). When the frequency matches the interpulse spacing, NV1-3 show positive pair-wise correlations from the correlated phase they accumulate due to the shared noise, and NV4 shows negative pair-wise correlations with NV1-3 because of the opposite phase accumulation.
Our work opens the door to high-throughput measurements of magnetic fields across large areas with nanoscale precision, enabling multiplexed relaxometry or noise spectroscopy of inhomogenous material systems with nanoscale resolution [14, 25, 43], as well as high-throughput diamond surface characterization and nanoscale NMR measurements [35, 40, 31]. Additionally, our multiplexing platform can be readily extended to accomplish low readout noise measurements of many more NV centers. Provided 10 W of laser power, we could address 500 NV centers resulting in more than $10^{5}$ measured two-point correlators. Massively multiplexed covariance magnetometry will offer a unique probe of correlated phenomena in condensed matter systems [32]; access to the intrinsic length- and time-scales of magnetic field dynamics with nanoscale precision would enable study of the rich noise dynamics found for instance in hydrodynamic electron flow [1, 23, 42], magnetic 2D materials [43, 25], and thin-film superconductors [11]. Furthermore, multiplexed super-resolution covariance magnetometry [33] could report on many two-point correlators with nanometer separations simultaneously, speeding up measurements of systems where dynamics are expected to have correlation length scales below the diffraction limit, such as spin liquids [9].

### v2

@@
Kai-Hung Cheng Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Zeeshawn Kazi Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Jared Rovny Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Bichen Zhang Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Lila S. Nassar Princeton University, Department of Physics, Princeton, NJ 08544, USA    Jeff D. Thompson Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA    Nathalie P. de Leon Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ 08544, USA
(August 25, 2024)
@@
Single nitrogen vacancy (NV) centers in diamond have been used extensively for high-sensitivity nanoscale sensing, but conventional approaches use confocal microscopy to measure individual centers sequentially [taylor_high-sensitivity_2008, maze_nanoscale_2008, lovchinsky_nuclear_2016, kucsko_nanometre-scale_2013, kolkowitz_sensing_2012, dolde_electric-field_2011], limiting throughput and access to non-local physical properties. Here we design and implement a multiplexed NV sensing platform that allows us to read out many single NV centers simultaneously using a low-noise camera. Using this platform, we coherently manipulate and read out the spin states of hundreds of individual NV centers in parallel, achieving comparable magnetic field sensitivity to confocal measurements. We also implement a parallelized version of spin-to-charge-conversion readout for low NV center spin state readout noise [shields_efficient_2015] and use it to demonstrate multiplexed covariance magnetometry [rovny_nanoscale_2022], in which we measure six two-point magnetic field correlators from four NV centers simultaneously. The number of correlators we can measure is limited only by the available laser power, opening the door to massively multiplexed covariance magnetometry. In contrast to scanning probes for imaging spatially varying magnetic fields [marchiori_nanoscale_2022], our approach measures the temporal dynamics of the magnetic field at many precisely defined positions simultaneously, from which the local noise spectrum and non-local properties such as correlation functions can be computed. This will find immediate applications to studying condensed matter phenomena that are characterized by the noise spectrum or correlation functions [rovny_new_2024], including quantum phase transitions [machado_quantum_2023], dynamics far from equilibrium [Xue2024, andersen_electron-phonon_2019, zhang_nanoscale_2024], magnetic order [ziffer_quantum_2024, Li2024a], and correlated electron phenomena [Chatterjee2019].
@@
Nitrogen vacancy (NV) centers in diamond have long spin coherence times across a wide range of temperatures combined with optical spin initialization and readout, enabling broad applications in nanoscale sensing of quantities such as magnetic field, electric field, and temperature [taylor_high-sensitivity_2008, maze_nanoscale_2008, lovchinsky_nuclear_2016, kucsko_nanometre-scale_2013, kolkowitz_sensing_2012, dolde_electric-field_2011]. NV centers are typically deployed either as individual sensors in bulk diamond [taylor_high-sensitivity_2008, maze_nanoscale_2008, lovchinsky_nuclear_2016, sutula_large-scale_2023, kolkowitz_sensing_2012, Childress2006], scanning tip geometries [Maletinsky2012, ariyaratne_nanoscale_2018], or in dense ensembles of NV centers that are not individually resolvable [glenn_high-resolution_2018, glenn_micrometer-scale_2017, barry_optical_2016, kazi_direct_2024, rodgers_diamond_2024]. Resolvable NV centers offer nanometer sensing volumes [Pham2016] and high sensitivity even for NV centers within a few nanometers of the surface with proper surface preparation [sangtawesin_origins_2019, yuan_charge_2020] but suffer from slow measurements due to the high readout noise of the NV center spin state [doherty_nitrogen-vacancy_2013, Barry2020]. NV center ensembles offer fast, wide-field measurements by averaging over the signal measured by many NV centers [glenn_high-resolution_2018, glenn_micrometer-scale_2017, Barry2020, kazi_wide-field_2021] at the cost of micron-scale spatial averaging imposed by the diffraction limit.
@@
In this work, we demonstrate a multiplexed NV sensing platform that combines the best of these modalities, realizing parallel, wide-field measurements of many individually resolvable NV centers using a low-noise camera. We first demonstrate parallel continuous-wave (CW) ODMR of hundreds of resolvable NV centers, achieving a magnetic field sensitivity within a factor of 5 of confocal measurements of single centers in the same sample. We extend this technique to spatially resolved spin measurements, performing parallel relaxometry ($T_{1}$ measurements) of 70 NV centers. Next, we demonstrate enhanced parallel spin readout by implementing spin-to-charge conversion (SCC) using a spatial light modulator (SLM) to focus a high-power ionization laser onto multiple NV centers simultaneously. We realize efficient ionization and multiplexed single-shot charge state readout with a fidelity of 88%, which enables multiplexed SCC spin readout with low readout noise, around ten times the quantum projection noise limit. This architecture allows us to measure magnetic field correlations between four individually resolved NV centers in parallel. The scalability of the multiplexed SCC is limited by the ionization laser power, and we expect that these techniques can be extended to hundreds of NV centers in future work, enabling parallel nanoscale sensing of non-local properties at many length scales [rovny_new_2024].
@@
Our multiplexing platform (Fig. 1a) uses an electron-multiplying charge-coupled device (EMCCD) camera to detect NV center fluorescence and a microwave printed-circuit board in front of the objective to deliver microwave excitation [yuan_instructional_2024]. Using CW wide-field green (532 nm) excitation, we image several hundred NV centers with approximately 2 seconds of integration time on the camera (Fig. 1b). We demonstrate parallel spin control and readout by measuring the optically-detected magnetic resonance (ODMR) spectrum of all the NV centers in the field of view. We simultaneously measure several hundred ODMR spectra in parallel (Fig. 1d) by recording the fluorescence emitted by each NV center under CW excitation as a function of applied microwave frequency and normalizing by a reference measurement with no applied microwaves (Fig. 2a). NV centers exist in four possible crystallographic orientations [taylor_high-sensitivity_2008], and three distinct orientations are observed here (Fig. 1d, different colors) because two are degenerate under the static magnetic field alignment. From the ODMR spectra, we calculate the DC magnetic field sensitivity of each individual NV center and find the median DC sensitivity is 24 µT Hz-1/2 (Fig. 2b) compared to a median sensitivity of 5.5 µT Hz-1/2 obtained with confocal measurements of the same diamond sample (Extended Data Figs. 1 and 2).
@@
Parallel readout also enables the multiplexed measurement of NV center dynamics. Pulsed measurements are challenging because the EMCCD camera data transfer time of 10 ms is much longer than the typical readout time for NV center spin contrast of 300 ns, determined by the lifetime of the shelving state [doherty_nitrogen-vacancy_2013]. For wide-field pulsed measurements using the camera, we repeat the experiment multiple times while integrating the fluorescence on the camera (Fig. 2d[kazi_wide-field_2021]. This reduces the overhead of frame transfer time and suppresses the noise from camera readout but lowers the fluorescence contrast between the NV spin states because the camera also records fluorescence during the spin reinitialization between experiments. With this approach, we are able to measure Rabi oscillations of many NV centers simultaneously (Fig. 2c). We demonstrate the application of parallel pulsed experiments to measure the spin relaxation time $T_{1}$ of over 70 NV centers in parallel (Fig. 2e). The measured $T_{1}$ times in this sample, which has NV centers implanted within approximately 10 nm of the surface, range from a few hundred microseconds to over 4 ms, consistent with past measurements of similarly prepared samples [sangtawesin_origins_2019, Bluvstein2019].
@@
We next demonstrate multiplexed low noise readout of the NV center spin state using SCC, which is critical for covariance magnetometry [rovny_nanoscale_2022] as the time required to measure a magnetic field correlation signal with a given signal-to-noise ratio scales as $\sigma_{\text{R}}^{4}$ where $\sigma_{\text{R}}$ is defined as the single-shot spin state readout noise [Barry2020]. SCC achieves lower $\sigma_{\text{R}}$ compared to green readout by using a high-power laser pulse to selectively ionize the NV center based on its spin state, followed by a high-fidelity charge state readout [shields_efficient_2015]. The selective ionization step typically requires $\sim 10$ mW of optical power per NV center, so implementing SCC with a wide-field laser beam would require $>$100 W of power for a 20 µm $\times$ 20 µm region. Thus, to realize multiplexed SCC in a power efficient manner, we use an SLM to create a programmable pattern of laser beams that only excite targeted NV centers (Fig. 1c). For simplicity, we use the same illumination pattern and wavelength (orange, 594 nm) for both ionization and charge state readout.
@@
We implement multiplexed charge state readout with the SLM and the camera through the following steps, outlined in Fig. 3a-c (Methods Section I). First, we take an image of all the NV centers using wide-field green excitation and identify the NV centers that we want to target (Fig. 3a). Second, we run a weighted Gerchberg-Saxton algorithm to generate a phase pattern on the SLM to create the laser spots at the fitted NV center locations (Fig. 3b[nogrette_single-atom_2014]. To read out the charge state, we excite the NV centers with a low-power orange laser and measure the photon counts per exposure by thresholding individual pixels in a region around each NV center and then summing the thresholded counts (Fig. 3c). In this low-photon regime, we binarize the image such that each pixel measures either 0 or 1 photons and do not account for higher per-pixel counts, which allows us to take full advantage of the camera’s readout statistics [basden_photon_2003]. A representative histogram of photon counts per camera frame is shown in Fig. 3d (Extended Data Fig. 3), revealing two peaks that correspond to the NV0 (dark) and NV- (bright) charge states under orange illumination. The measured charge state readout fidelity using the camera (88.3%) is similar to that obtained using confocal readout of the same sample (86.3%) with the same objective lens and readout parameters (Extended Data Fig. 4).
@@
Having achieved multiplexed charge-state measurements using targeted excitation, we then add a spin-selective ionization pulse to implement multiplexed SCC by modulating the power of the orange laser between ionization ($\sim$10 mW per spot) and readout ($\sim$20 µW per spot). We compare the charge state readout statistics for a representative NV center spin prepared in either the $m_{s}=0$ or $m_{s}=-1$ spin state to benchmark $\sigma_{\text{R}}$ (Fig. 3e). The spin readout noise for the data shown in Fig. 3e is $\sigma_{\text{R}}\approx 12$, a four-fold improvement over green readout in a confocal microscope [Barry2020] and a ten-fold improvement over wide-field green readout (Extended Data Fig. 5). However, our scheme performs worse than previous implementations of SCC using a red (637 nm) laser for ionization ($\sigma_{\text{R}}\sim 5$[rovny_nanoscale_2022], which we attribute to a higher charge state recombination rate from the orange laser [aslam_photo-induced_2013], in agreement with a rate equation model [wirtitsch_exploiting_2023] (Extended Data Fig. 6).
@@
Equipped with multiplexed SCC, we demonstrate parallelized covariance magnetometry by reconstructing non-local magnetic field correlators $\braket{B(\vec{x}_{i})B(\vec{x}_{j})}$ from simultaneous measurements of NV centers at different locations $\vec{x}_{i}$ [rovny_nanoscale_2022]. We target 4 NV centers (Fig. 4a) from which we simultaneously measure (${4\choose 2}=6$ two-point correlators. To demonstrate that the measured correlations result from spin dynamics rather than spurious correlations [rovny_nanoscale_2022], we use a separate green laser to prepare one of the NV centers (NV4) in the opposite spin state from the other three (NV1-3) (Fig. 4a,b). To verify our ability to detect two-point correlators with our achievable readout noise, we first measure correlations introduced by direct NV spin driving rather than noise sensing (Fig. 4b). This allows us to rapidly drive the NV centers to mitigate the effect of decoherence under free evolution, as well as to circumvent quantum projection noise as we are able to intentionally drive each NV center between pure states $m_{s}=0$ and $m_{s}=-1$. In this driven-spin correlation experiment, we interleave two pulse sequences: in the even-numbered experiments, we simultaneously drive all the NV centers around the Bloch sphere by an angle $\theta$, while in the odd-numbered experiments we drive the NV centers by an angle $\theta+\pi$ (Fig. 4b). At the end of each experiment we read out a single camera frame, simultaneously measuring the four signals $s_{i}$ where $i=1...4$ indexes the NV center. Lastly, we use the set of acquired signals across experiments $S_{i}=\{s_{i,k}\}$, where $k$ indexes the experiment number, to compute the Pearson correlation for each NV pair: $r_{ij}=\text{Cov}(S_{i},S_{j})/(\sigma_{i}\sigma_{j})$, where Cov is the covariance and $\sigma_{i}$ is the standard deviation of $S_{i}$. From the computed Pearson correlations, we subtract off a uniform background correlation measured using an independent calibration experiment (Extended Data Fig. 7); the resulting correlations are shown in Fig. 4c. As expected for each pair of NV centers prepared in the same initial spin state (NV1-3), the correlations between the measured signals are maximized when $\theta=n\pi$ for integer $n$, as the two spins flip synchronously between $m_{s}$ = 00 and $-1$, and are minimized for $\theta=(n+1/2)\pi$, since the readout projects the two spins randomly and independently onto $m_{s}$ = 00 or $-1$. Conversely, each NV center’s correlation with NV4 is negative, as NV4 has been prepared in the opposite orientation and thus acquires an opposite phase. In addition, the amplitude of the correlation oscillation indicates the maximum measurable correlation, which is $1/\sigma_{R_{i}}\sigma_{R_{j}}$. From Fig. 4c, we estimate $\sigma_{R_{i}}=$ $15\pm 1$, consistent with the data in Fig. 3f.
@@
We also use multiplexed covariance magnetometry to perform covariance noise spectroscopy of an applied AC magnetic field [rovny_nanoscale_2022]. As before, we prepare NV1-3 in the same spin state and prepare NV4 in the opposite state. We then use an XY8 sensing sequence [Meiboom1958] with a fixed interpulse spacing, and sweep the frequency of the applied AC signal (Fig. 4d). When the frequency matches the interpulse spacing, NV1-3 show positive pair-wise correlations from the correlated phase they accumulate due to the shared noise, and NV4 shows negative pair-wise correlations with NV1-3 because of the opposite phase accumulation.
The platform demonstrated here can be readily extended to accomplish low readout noise measurements of many more NV centers. Provided 10 W of ionization laser power, we could address 500 NV centers resulting in more than $10^{5}$ measured two-point correlators. Moreover, performing ionization at a more optimal wavelength is expected to reduce $\sigma_{\text{R}}$ further. Finally, extending these techniques to measure correlations below the diffraction limit [Rovny2024b] would allow for the measurement of many two-point correlators with nanometer separations simultaneously.
Our platform opens the door to high-throughput measurements of magnetic fields across large areas with nanoscale localization, with immediate applications in magnetometry of condensed matter dynamics. Just as scanning NV magnetometry has advanced our understanding of magnetic ordering [Thiel2019], transport [Ku2020, Jenkins2022, Vool2021], and superconductivity [Thiel2016, Chen2024] through high-resolution imaging of magnetic fields, our platform will provide a new window into dynamics through wide-field mapping of local noise spectra and non-local correlation functions at many length scales simultaneously. Specifically, local noise measurements can provide a quantitative probe of dynamics [kolkowitz_probing_2015, dwyer_probing_2022], for instance in measuring critical exponents of phase transitions in 2D antiferromagnets that have vanishing stray fields [machado_quantum_2023, ziffer_quantum_2024]. Additionally, access to non-local correlation functions with nanoscale precision will enable study of the dynamics in systems that are fundamentally characterized by correlations, such as in hydrodynamic electron flow [Agarwal2017, Lucas2018, zhang_nanoscale_2024], magnetic 2D materials [ziffer_quantum_2024, machado_quantum_2023], vortex dynamics in superconductors [curtis_probing_2024, chen_when_2024], and spin liquids [Chatterjee2019, Broholm2020].